Nnesha (nnesha):

A small business owner made $40,000 the first year he owned his store and made an additional 7% over the previous year in each subsequent year. Find how much he made during his fourth year of business. Find his total earnings during the first four years.

OpenStudy (rational):

we may use \(n\)th term of geometric sequence formula to find the amout he makes during fourth year : \[\large a_n = a_1r^{n-1}\]

OpenStudy (rational):

here \(r = 1.07\) and \(a_1 = 40,000\)

Nnesha (nnesha):

don't we have to use this one ?? \[\rm A =P(1+\frac{ r }{ n })^{nt}\] or A = p(1+r)^t

OpenStudy (rational):

that also works!

OpenStudy (rational):

otherwise work each term manually as there are only four terms

OpenStudy (rational):

first year : 40,000 second year : 40,000 + (7% of 40,000) = ?

Nnesha (nnesha):

\[\huge\rm a_4= 40000(1.07)^{4-1}\] 49,001.72

OpenStudy (rational):

that looks good!

Nnesha (nnesha):

what about this one "A small business owner made \(\large\rm{$3000}\) the first year he owned his store and made an additional \[14%\] over the previous year in each subsequent year. Find how much he made during his fourth year of business. Find his total earnings during the first four years. \[a_4=30000(1.14)^{3}\] =4444.32 is it right ??

OpenStudy (rational):

this looks more like arithmetic sequence right ? because the growth amount is same each year : $14

Nnesha (nnesha):

14%**

OpenStudy (rational):

then you're correct !

Nnesha (nnesha):

but answer choices are a) $33,745.92 , 127,393.92 <----------------A is right answer :o b) $82,230.00 ; $213,120.00 c)$1.92; $31249.92 d)$115248.00:$328,368.00

OpenStudy (rational):

is the first year amount $3,000 or $30,000 ?

Nnesha (nnesha):

$30,000

OpenStudy (rational):

Okay what about the growth factor ? is it 14% or just 4% ?

Nnesha (nnesha):

_-_ 4% ;p

Nnesha (nnesha):

thanks!!

OpenStudy (rational):

np :)